# Genesis: Multimodal Driving Scene Generation with Spatio-Temporal and Cross-Modal Consistency

### v3

@@ Abstract
We present Genesis, a unified framework for joint generation of multi-view driving videos and LiDAR sequences with spatio-temporal and cross-modal consistency. Genesis employs a two-stage architecture that integrates a DiT-based video diffusion model with 3D-VAE encoding, and a BEV-aware LiDAR generator with NeRF-based rendering and adaptive sampling. Both modalities are directly coupled through a shared latent space, enabling coherent evolution across visual and geometric domains. To guide the generation with structured semantics, we introduce DataCrafter, a captioning module built on vision-language models that provides scene-level and instance-level supervision. Extensive experiments on the nuScenes benchmark demonstrate that Genesis achieves state-of-the-art performance across video and LiDAR metrics (FVD 16.95, FID 4.24, Chamfer 0.611), and benefits downstream tasks including segmentation and 3D detection, validating the semantic fidelity and practical utility of the generated data.
@@ 1 Introduction
As shown in Fig. 1, existing driving scene generation methods typically focus on generating data in a single modality, usually RGB videos bevgen ; magicdrive ; drivedreamer ; drivedreamer2 ; drivewm ; panacea ; dive ; ma2024unleashing or LiDAR point clouds lidargen ; lidardm ; transfusion ; cameralidar ; deepfusion ; bevfusion . While these approaches have significantly advanced the field of driving scene generation, they overlook the synergistic potential of multimodal generation and lack consistency in aligning RGB videos with various sensor data, leading to limitations in real-world applications. Many of these methods rely on one-step layout-to-data pipelines conditioned solely on coarse spatial priors, such as BEV maps street ; magicdrive ; tune ; bevcontrol ; drivedreamer ; drivedreamer2 ; drivewm or 3D boxes kitti ; nuscenes ; waymo , which restricts their ability to capture complex scene dynamics and fine-grained semantics. To propel driving scene generation into the multimodal domain, recent work, such as UniScene li2024uniscene introduces occupancy grids for multimodal generation, but its decoupled design and weak semantic supervision limit cross-modal alignment and scene fidelity. Moreover, like most existing multimodal generation approaches, current methods generally rely on limited semantic supervision—typically in the form of coarse labels or generic captioning models—without fully leveraging the fine-grained descriptive capabilities of modern vision-language models (VLMs) internvl ; liu2023visual ; qwen . This lack of structured semantic grounding restricts the fidelity, controllability, and contextual alignment of generated scenes.
Therefore, advancing multimodal driving scene generation requires more than unimodal fidelity—it demands tight cross-modal alignment and fine-grained semantic grounding to ensure spatiotemporal coherence. However, existing methods suffer from three core limitations: (1) they often decouple video and LiDAR generation, weakening cross-modal consistency; (2) they rely on intermediate representations like occupancy grids that may cause information loss; (3) they provide limited semantic supervision, typically in the form of coarse layout maps or generic captions, which hinders scene-level controllability and realism.
@@
  • Unified Multimodal Generation Architecture. Genesis employs a unified pipeline where both video and LiDAR branches operate within a shared latent space. Vision and geometry are directly coupled through a novel cross-modal conditioning mechanism, enabling consistent temporal evolution and geometric alignment across modalities without relying on occupancy or voxel intermediates.
  • Structured Semantic Supervision via DataCrafter. To improve semantic controllability, we introduce DataCrafter, a structured captioning module built upon vision-language models. It extracts multi-view, scene-level and instance-level descriptions that are fused into dense, language-guided priors. These captions provide detailed semantic guidance to both video and LiDAR generators, resulting in outputs that are not only realistic but also interpretable and controllable.
@@ 3 Method
Overview. As illustrated in Fig. 2, Genesis adopts a unified two-branch architecture to jointly synthesize multimodal driving scenes, including videos and LiDAR point clouds. The video generation branch leverages a DiT-based spatiotemporal diffusion backbone with a 3D-VAE encoder to capture fine-grained visual dynamics (Sec. 3.2), while the LiDAR branch employs a BEV-aware AE with NeRF-style rendering and adaptive sampling to ensure accurate geometric reconstruction (Sec. 3.3). Both branches are jointly conditioned on road sketches, intermediate video latent features, and structured scene semantics to maintain strong cross-modal consistency. Semantic priors such as captions are derived from our DataCrafter module (Sec. 3.1) and injected as global conditioning to enhance alignment across modalities. More architectural details are provided in the supplementary material (Sec. A).
@@ 3.2 Video Generation Model
Figure 5: Illustration of the BEV sketch projection pipeline, which enables view-consistent and semantically grounded video generation.
@@
To ensure spatiotemporal consistency of control signals, a 3D Variational Autoencoder (VAE) is employed to compress the multi-frame BEV sketch $\mathcal{S}^{l}$ into a latent representation $z\in\mathbb{R}^{f\times c\times h\times w}$. The decoder reconstructs the BEV semantics from the denoised tokens. The training objective is defined as:
$\mathcal{L}_{\text{sem}}=\mathcal{L}_{\text{CE}}+\alpha\mathcal{L}_{\text{KL}}+\beta\mathcal{L}_{\text{Lovasz}},$ (5)
where $\mathcal{L}_{\text{CE}}$ denotes cross-entropy loss, $\mathcal{L}_{\text{KL}}$ is the KL divergence term, and $\mathcal{L}_{\text{Lovasz}}$ encourages structured boundary-aware predictions.
@@
$e_{\text{cap}}=\mathcal{E}_{\text{text}}(\text{caption}),\quad z_{s}=\mathcal{E}_{\text{sketch}}(s).$ (6)
@@ 3.3 Lidar Generation Model
Spatiotemporal Diffusion with ControlNet-Branch. With the latent embedding from the Point Cloud AutoEncoder Module, we design a diffusion network to learn this representation to generate realistic point clouds. Similar to the structure adopted in the video generation, we use a dual DiT-based network with ControlNet assistance. For semantic condition input, we use scene captions and road sketches to provide structured scene priors. To keep cross modal consistency, we use the camera branch’s RGB video outputs as another condition input. We adopt widely used LSS philion2020lift algorithm to convert the perspective images to BEV feature. For geometric condition input, we encode 3D bounding boxes into embeddings for providing accurate foreground information. The image BEV feature is concatenated with the road sketch latent feature and then send into the ControlNet branch as input. For a given latent token $z_{i}$ at block $i$, the cross-attention operation cloud be formulated as:
@@ 4.2 Qualitative comparison and Versatile Generation Ability
Figure 6: Qualitative comparison of video generation. From top to bottom: (1) Ground-truth images, (2) Road sketch input, (3) Panacea panacea , (4) MagicDrive magicdrive , (5) Ours. Panacea suffers from hallucinated textures and geometric misalignment. MagicDrive shows vehicle distortion and broken structures. In contrast, ours preserves accurate layout, object shapes, and background integrity.
@@ 4.6 Ablation Studies
We conduct an ablation study to assess the effectiveness of the proposed DataCrafter and PoseNet modules, which introduce structured semantic supervision into the video generation pipeline. As shown in Tab. 3, removing DataCrafter causes a notable degradation in generation quality, with FVDmulti increasing from 85.91 to 117.49 and FIDmulti rising from 15.20 to 22.32. Including the PoseNet module further improves performance, reducing FVDmulti to 83.10 and FIDmulti to 14.90. These results underscore the importance of semantic guidance for maintaining temporal consistency and structural fidelity. As illustrated in Fig. 8, caption-conditioned generation produces sharper structures and more coherent layouts, whereas the ablated variant exhibits missing objects and visual distortions. In addition, PoseNet contributes fine-grained dynamic cues by localizing human anatomical keypoints.
@@ 5 Conclusion
We propose Genesis, a unified framework for the joint generation of multi-view videos and LiDAR point clouds in autonomous driving scenarios. By integrating structured semantic priors through the DataCrafter module and enforcing cross-modal alignment via a latent conditioning pipeline, Genesis effectively bridges the gap between visual and geometric modalities. This design enables the synthesis of high-fidelity, spatiotemporally coherent, and semantically faithful multimodal sequences. Extensive experiments across a range of benchmarks demonstrate that Genesis achieves sota performance in both video and LiDAR generation tasks, highlighting its robustness and scalability.

### v4

@@ Abstract
We present Genesis, a unified framework for joint generation of multi-view driving videos and LiDAR sequences with spatio-temporal and cross-modal consistency. Genesis employs a two-stage architecture that integrates a DiT-based video diffusion model with 3D-VAE encoding, and a BEV-represented LiDAR generator with NeRF-based rendering and adaptive sampling. Both modalities are directly coupled through a shared condition input, enabling coherent evolution across visual and geometric domains. To guide the generation with structured semantics, we introduce DataCrafter, a captioning module built on vision-language models that provides scene-level and instance-level captions. Extensive experiments on the nuScenes benchmark demonstrate that Genesis achieves state-of-the-art performance across video and LiDAR metrics (FVD 16.95, FID 4.24, Chamfer 0.611), and benefits downstream tasks including segmentation and 3D detection, validating the semantic fidelity and practical utility of the synthetical data.
@@ 1 Introduction
As shown in Fig. 1, existing driving scene generation methods typically focus on generating data in a single modality, usually RGB videos bevgen ; magicdrive ; drivedreamer ; drivedreamer2 ; drivewm ; panacea ; dive ; ma2024unleashing or LiDAR point clouds lidargen ; lidardm ; transfusion ; cameralidar ; deepfusion ; bevfusion . While these approaches have significantly advanced the field of driving scene generation, they overlook the synergistic potential of multimodal generation and lack consistency in aligning RGB videos with various sensor data, leading to limitations in real-world applications. Many of these methods rely on one-step layout-to-data pipelines conditioned solely on coarse spatial priors, such as BEV maps street ; magicdrive ; tune ; bevcontrol ; drivedreamer ; drivedreamer2 ; drivewm or 3D boxes kitti ; nuscenes ; waymo , which restricts their ability to capture complex scene dynamics and fine-grained semantics. To propel driving scene generation into the multimodal domain, UniScene li2024uniscene involves occupancy grids in for multimodal generation, while the acquisition of occupancy labels is highly costly, which severely limits its application. Moreover, most existing generation approaches generally rely on limited semantic annotations, typically in the form of coarse labels or generic captioning models, without fully leveraging the fine-grained descriptive capabilities of modern vision-language models (VLMs) internvl ; liu2023visual ; qwen . This lack of structured semantic grounding restricts the fidelity, controllability, and contextual alignment of generated scenes.
Therefore, advancing multimodal driving scene generation requires more than unimodal fidelity—it demands tight cross-modal alignment and fine-grained semantic grounding to ensure spatio-temporal coherence. However, existing methods suffer from three core limitations: (1) they often decouple video and LiDAR generation, weakening cross-modal consistency; (2) they rely on intermediate representations like occupancy grids that may cause information loss; (3) they provide limited semantic supervision, typically in the form of coarse layout maps or generic captions, which hinders scene-level controllability and realism.
@@
  • Unified Multimodal Generation Architecture. Genesis employs a unified pipeline where both video and LiDAR branches use shared conditional inputs, including scene descriptions and scene layouts, etc. To further guarantee the consistency between point clouds and image, BEV features extracted from projected image features are incorporated as conditional inputs into the Lidar diffusion model. This approach enhances cross-modality consistency without relying on intermediaries like occupancy grids.
  • Structured Semantic Captions via DataCrafter. To improve semantic controllability, we introduce DataCrafter, a structured captioning module built upon vision-language models. It extracts multi-view, scene-level and instance-level descriptions that are fused into dense, language-guided priors. These captions provide detailed semantic guidance to both video and LiDAR generators, resulting in outputs that are not only realistic but also interpretable and controllable.
@@ 3 Method
Overview. As illustrated in Fig. 2, Genesis adopts a unified two-branch architecture to jointly synthesize multimodal driving scenes, including videos and LiDAR point clouds. The video generation branch leverages a DiT-based spatiotemporal diffusion backbone with a 3D-VAE encoder to capture fine-grained visual dynamics (Sec. 3.2), while the LiDAR branch employs a BEV-aware AE with NeRF-style rendering and adaptive sampling to ensure accurate geometric reconstruction (Sec. 3.3). Both branches are jointly conditioned on scene layouts, intermediate video latent features, and structured scene semantics to maintain strong cross-modal consistency. Semantic priors such as captions are derived from our DataCrafter module (Sec. 3.1) and injected as global conditioning to enhance alignment across modalities. More architectural details are provided in the supplementary material (Sec. A).
@@ 3.2 Video Generation Model
Figure 5: Illustration of the scene layouts projection pipeline, which enables view-consistent and semantically grounded video generation.
@@
To ensure spatiotemporal consistency of control signals, a 3D Variational Autoencoder (VAE) is employed to compress the multi-frame scene layouts $\mathcal{S}^{l}$ and images into a latent representation $z\in\mathbb{R}^{f\times c\times h\times w}$. The decoder reconstructs the RGB images from the denoised latents.
@@
$e_{\text{cap}}=\mathcal{E}_{\text{text}}(\text{caption}),\quad z_{s}=\mathcal{E}_{\text{layout}}(s).$ (5)
@@ 3.3 Lidar Generation Model
Spatiotemporal Diffusion with ControlNet-Branch. With the latent embedding from the Point Cloud AutoEncoder Module, we design a diffusion network to learn this representation to generate realistic point clouds. Similar to the structure adopted in the video generation, we use a dual DiT-based network with ControlNet assistance. For semantic condition input, we use scene captions and scene layouts to provide structured scene priors. To keep cross modal consistency, we use the camera branch’s RGB video outputs as another condition input. We adopt widely used LSS philion2020lift algorithm to convert the perspective images to BEV feature. For geometric condition input, we encode 3D bounding boxes into embeddings for providing accurate foreground information. The image BEV feature is concatenated with the scene layouts latent feature and then send into the ControlNet branch as input. For a given latent token $z_{i}$ at block $i$, the cross-attention operation cloud be formulated as:
@@ 4.2 Qualitative comparison and Versatile Generation Ability
Figure 6: Qualitative comparison of video generation. From top to bottom: (1) Ground-truth images, (2) Scene Layouts input, (3) Panacea panacea , (4) MagicDrive magicdrive , (5) Ours. Panacea suffers from hallucinated textures and geometric misalignment. MagicDrive shows vehicle distortion and broken structures. In contrast, ours preserves accurate layout, object shapes, and background integrity.
@@ 4.6 Ablation Studies
We conduct an ablation study to assess the effectiveness of the proposed DataCrafter and PoseNet modules, which introduce structured semantic captions into the video generation pipeline. As shown in Tab. 3, removing DataCrafter causes a notable degradation in generation quality, with FVDmulti increasing from 85.91 to 117.49 and FIDmulti rising from 15.20 to 22.32. Including the PoseNet module further improves performance, reducing FVDmulti to 83.10 and FIDmulti to 14.90. These results underscore the importance of semantic guidance for maintaining temporal consistency and structural fidelity. As illustrated in Fig. 8, caption-conditioned generation produces sharper structures and more coherent layouts, whereas the ablated variant exhibits missing objects and visual distortions. In addition, PoseNet contributes fine-grained dynamic cues by localizing human anatomical keypoints.
@@ 5 Conclusion
We propose Genesis, a unified framework for joint multi-view video and LiDAR point cloud generation in autonomous driving. By integrating structured semantic priors via the DataCrafter module and enforcing cross-modal alignment through a shared conditioning pipeline, Genesis bridges visual and geometric modalities. This enables high-fidelity, spatiotemporally coherent, and semantically consistent multimodal sequence synthesis. Experiments on nuScenes show Genesis achieves SOTA performance in both video and LiDAR generation, demonstrating its robustness and scalability.